# 03F40 (primary), 03D20

### v2

@@ 5 The Ouroboros Dynamical Hierarchy
We have shown that fractal Goodstein processes terminate, regardless of $\mathcal{B}$; in view of Example 2.12, this can be seen as an alternative proof of Goodstein’s original theorem if we set $\mathcal{B}_{i}=\{i+2\}$. However, for other choices of $\mathcal{B}$, the proof-theoretic strength of termination is much higher. Our goal is to define a dynamical hierarchy which will yield a Goodstein principle independent of $\sf KP$. This will be achieved by defining some specially tailored good successors $B_{+i}$ for a given base hierarchy $B$, where $i$ is a natural number parameter. Our dynamical hierarchy, which will be denoted $\mathcal{O}$, will then be defined so that $\mathcal{O}_{i+1}=(\mathcal{O}_{i})_{+i}$ for all $i\in\mathbb{N}$. The initial base hierarchy $\mathcal{O}_{0}$ is not too important and we may simply let $\mathcal{O}_{0}=\{2\}$, mirroring Goodstein’s original theorem.
Fix a base hierarchy $B$; let us sketch how $B_{+i}$ is to be constructed. We hinted in the introduction about the idea of changing the base of $n\in\mathbb{N}$ to $n$ itself when performing the upgrade operator. With the machinery we have developed, this would amount to finding a good successor $C$ for $B$ such that, for $b={\rm Base}_{B}(n)$, ${\uparrow}n=\langle\begin{smallmatrix}n\\ b\end{smallmatrix}\rangle n$. Suppose that a hypothetical such $C$ is given and let all transformations be from $B$ to $C$.
As per Definition 2.5, replacing the base of $n$ by $n$ would require ${\uparrow}(n-1)<\langle\begin{smallmatrix}n\\ b\end{smallmatrix}\rangle n<S_{C}(n)$. The first inequality cannot always be guaranteed to hold, since ${\uparrow}(n-1)$ may be quite large. Instead, we can let ${\uparrow}n=\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n$, where $d$ itself is of the form $\langle\begin{smallmatrix}d^{\prime}\\ b\end{smallmatrix}\rangle n$. The number $d$ is not equal to $n$, but it is at least as large, and in view of properties such as Lemmas 3.82 and 4.18, can be regarded as a ‘copy’ of $n$.
We need to have $d^{\prime}\geq{\rm base}_{C}({\uparrow}(n-1))$, but aside from that, we have some liberty in choosing the value of $d^{\prime}$. In fact, we may even have $d^{\prime}=\langle\begin{smallmatrix}d^{\prime\prime}\\ b\end{smallmatrix}\rangle n$, and we can iterate this idea any finite number of steps. Thus the general picture will be that, for a number $n$, we will add a sequence of bases $d_{1},\ldots,d_{i}$, with $d_{j+1}=\langle\begin{smallmatrix}d_{j}\\ b\end{smallmatrix}\rangle n$. The restriction $d_{j}\geq{\rm base}_{C}({\uparrow}(n-1))$ is obtained by letting $d_{0}$ be the largest base previously available, then adding this as an initial value to the sequence. Note that the value of $i$ can vary; if we follow this construction for a given $i$, we obtain $B_{+i}:=C$.
Since our goal is to show that Theorem 2.16 is not provable in $\sf KP$, the dynamical hierarchy we will construct will be a sort of ‘worst case scenario,’ so we want to add the new bases $\{d_{j}\}^{i}_{j=1}$ whenever possible. Nevertheless, this cannot be done for all $n\in\mathbb{N}$. If $n<\min B$, then $n$ must be fixed by the upgrade, and $\min B$ must be mapped to $\min C\geq\min B$. If we want $C$ to be a good successor of $B$, then whenever $\in B, n{\uparrow}n$ is determined by ${\uparrow}(n-1)$; namely, it must be the next multiple of ${\rm base}_{C}({\uparrow}(n-1))$. Likewise, if $\min B<n\notin B$ and $b\nmid n$, then for any $d\geq{\uparrow}b$, we have that $d\nmid\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n$ by Lemma 3.81, so $\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n$ cannot be a base if $d\in C$. Thus we will add extra bases only in the ‘critical’ case, where $\min B<n$ and ${\rm Base}_{B}(n)\mid n\notin B$. With this, arrive at the following construction.
Definition 5.1.
Let $B$ be a base hierarchy and $i\in\mathbb{N}\setminus\{0\}$. Say that $n>0$ is $B$-critical if ${\rm Base}_{B}(n)\mid n\notin B$. Let $\partial B$ denote the set of $B$-critical numbers.
We define a new base hierarchy $B_{+i}=\bigcup_{n<\infty}B_{+i}^{n}$, where $B_{+i}^{n}$ is defined recursively as follows. If $n\leq\min B$, define $B_{+i}^{n}=\{\min B+1\}$. For $n>\min B$, assume inductively that $B_{+i}^{m}$ has been defined and is finite for $m<n$, and let transformations be from $B$ to $B^{n-1}_{+i}$. Let $b={\rm base}_{B}(n)$ and $c=\max B_{+i}^{n-1}$.
  1. Case 1
    ($n\notin B\cup\partial B$). Set $B_{+i}^{n}=B_{+i}^{n-1}$.
  2. Case 2
    ($n\in B$). Let $a$ be least so that ${\uparrow}(n-1)<ca$, and set [^3] As a failsafe, we may set $C_{n}=C_{n-1}$ if ${\uparrow}(n-1)=\infty$; however, as we will see, this will never be the case. A similar remark applies to Case 3.
    $B_{+i}^{n}=B_{+i}^{n-1}\cup\big\{ca\}.$
  3. Case 3
    ($n\in\partial B$). Define $d_{j}=d_{j}(n)$ recursively by $d_{0}=c$ and
    $d_{j+1}=\langle\begin{smallmatrix}d_{j}\\ b\end{smallmatrix}\rangle n.$
    Then, set
    $B_{+i}^{n}=B_{+i}^{n-1}\cup\{d_{j}\}_{j=1}^{i}.$
We refer to a hierarchy of the form $B_{+i}$ as an ouroboros successor. We define the ouroboros dynamical hierarchy $\mathcal{O}=(\mathcal{O}_{i})_{i\in\mathbb{N}}$ to be the unique dynamical hierarchy with $\mathcal{O}_{0}=\{2\}$ and $\mathcal{O}_{i+1}=(\mathcal{O}_{i})_{+i}$.
The moniker is given since in the critical case, an upgrade of $n$ is used as the new base for $n$, as if it were ‘eating its own tail’.
Example 5.2.
In Example 2.10, we defined an operation $B\mapsto B_{+}$. This is in fact a special case of an ouroboros successor: specifically, $B_{+}=B_{+0}$. To see this, note that if $n$ is $B$-critical, then we are to add the bases $\{d_{j}(n)\}_{j=1}^{0}$; but this set is empty, hence $B^{n}_{+0}=B^{n-1}_{+0}$. It follows that only the clause for $n\in B$ produces new bases, and then it is easily checked to produce the same new base as in the definition of $B_{+}$.
Example 5.3.
Let $B=\{2\}$ and let us compute $B^{6}_{+1}$. Similarly to Example 2.10, $B^{3}_{+1}=\{3\}$, but now $B^{4}_{+1}=B^{3}_{+1}\cup\{d_{1}(4)\}$, so we do add a new base for $4$. We easily compute $d_{1}(4)=\langle\begin{smallmatrix}3\\ 2\end{smallmatrix}\rangle 4=\langle\begin{smallmatrix}3\\ 2\end{smallmatrix}\rangle 2^{2}=3^{3}$ and hence $B^{4}_{+1}=\{2,27\}$. Since $2\nmid 5\notin B$, $B^{5}_{+1}=B^{4}_{+1}$. On the other hand, $B^{6}_{+1}=B^{5}_{+1}\cup\{d_{1}(6)\}$, and since $27$ is the maximum element of $B^{5}_{+1}$, we have that $d_{1}(6)=\langle\begin{smallmatrix}27\\ 2\end{smallmatrix}\rangle 6=\langle\begin{smallmatrix}27\\ 2\end{smallmatrix}\rangle(2^{2}+2)=27^{27}+27$.
In general, we will add one new base for every even number, and thus $B_{+1}$ is infinite, even though $B$ has only one element.
Example 5.4.
Once again, let $B=\{2\}$ and let us compute $B^{4}_{+2}$. The construction is very similar, including the computation of $d_{1}(4)=27$. However, in this case, we have that $B^{4}_{+2}=B^{3}_{+2}\cup\{d_{1}(4),d_{2}(4)\}$. So, we must also compute $d_{2}(4)=\langle\begin{smallmatrix}27\\ 2\end{smallmatrix}\rangle 4=27^{27}$, and hence $B^{4}_{+2}=\{3,27,27^{27}\}$.
Finally, we remark that, while the value of $d_{1}(4)$ did not change compared to Example 5.3, that is no longer the case for $d_{1}(6)$. Instead, we have $d_{1}(6)=\langle\begin{smallmatrix}27^{27}\\ 2\end{smallmatrix}\rangle 6$, a number we will not even attempt to represent, let alone the gargantuan $d_{2}(6)=\langle\begin{smallmatrix}d_{1}(6)\\ 2\end{smallmatrix}\rangle 6$.
The takeaway is that $B^{n}_{+i}$ adds new bases only at ‘critical’ numbers, i.e. those satisfying $1<{\rm base}_{B}(n)\mid n\notin B$, in which case $i$ new bases are added. Since there are infinitely many such numbers in any base hierarchy, even a finite one, $B_{+i}$ is always infinite whenever $i>0$. Moreover, $d_{j}(n)$ depends also on $B$ and $i$, but we omit them from notation since they will always be fixed and clear from context.
It is useful to note that in our construction of $B_{+i}$, the new bases that are added are always larger than any previously available base, and indeed are multiples of these bases.
@@ 3. Case 3
Let $B$ be a base hierarchy and $i\in\mathbb{N}$. Let $C=B_{+i}$ and for each $n\in\mathbb{N}$, $C_{n}=B_{+i}^{n}$, $c_{n}=\max C_{n}$, ${\uparrow}_{n}={\uparrow}_{B}^{C_{n}}$, and $\langle\begin{smallmatrix}\cdot\\ \cdot\end{smallmatrix}\rangle_{n}=\langle\begin{smallmatrix}\cdot\\ \cdot\end{smallmatrix}\rangle_{B}^{C_{n}}$. Then, for every $n\in\mathbb{N}$ with $b={\rm Base}_{B}(n)$:
  1. 1.
@@
  4. 4.
    $C_{n}$ is a good successor for $B{\upharpoonright}_{n}$.
Proof 5.6.
@@
Proceed by induction on $n$. In this proof, ${\uparrow}={\uparrow}_{n-1}$ and $\langle\begin{smallmatrix}\cdot\\ \cdot\end{smallmatrix}\rangle=\langle\begin{smallmatrix}\cdot\\ \cdot\end{smallmatrix}\rangle_{n-1}$; note that ${\uparrow}x$ for $x<n$ and $\langle\begin{smallmatrix}y\\ x\end{smallmatrix}\rangle z$ when $x\leq n$ are finite by the induction hypothesis applied to 4 and Lemma 3.14.
@@
If $n\leq\min B$, then 4 is easy to check since $C_{n}=\{\min B+1\}$ and $\min B+1>\min B\geq 2$, so $C_{n}$ is a base hierarchy for $B{\upharpoonright}_{\min B}$. The other items are either also easy to check or hold vacuously. Similarly, if $C_{n}=C_{n-1}$, then all claims follow directly from the induction hypothesis, so we may assume $n>\min B$ and $C_{n}\neq C_{n-1}$. Consider the following cases.
  1. Case 1
    ($n\in B$). Let $b^{\prime}={\rm Base}_{B}(n-1)$. In this case, $C_{n}=C_{n-1}\cup\{c_{n}\}$ by definition. By the induction hypothesis applied to 2 and 3, $\langle\begin{smallmatrix}c_{n-1}\\ b^{\prime}\end{smallmatrix}\rangle(n-1)={\uparrow}_{n-1}(n-1)<c_{n}=S_{C}(c_{n-1})$, and since $\langle\begin{smallmatrix}c_{n-1}\\ b^{\prime}\end{smallmatrix}\rangle(n-1)\geq c_{n-1}$, we have that $c_{n}=S_{C_{n}}({\uparrow}(n-1))$, from which 1 follows. By definition, $c_{n}$ is a multiple of $c_{n-1}$, and if $c_{n-1}>c\in C_{n}$, then we may apply the induction hypothesis to $C_{n-1}$ to see that $c\geq 2$ and $c\mid S_{C_{n-1}}(c)=S_{C_{n}}(c)$. If $x<n$, then by Lemma 3.11, ${\uparrow}x\leq{\uparrow}(n-1)<c_{n}$, and since $C_{n-1}=C_{n}\upharpoonright_{c_{n}-1}$, 2 follows from Lemma 3.141. By Lemma 3.61, ${\uparrow}_{n}n=S_{C_{n}}({\uparrow}(n-1))=c_{n}=\langle\begin{smallmatrix}c_{n}\\ n\end{smallmatrix}\rangle_{n}n$, establishing 3. Since we chose $c_{n}$ to be the least multiple of $c_{n}-1$ above ${\uparrow}(n-1)$ and $C_{n-1}$ is a good successor for $B$ (and hence a base hierarchy), one readily checks that every element of $C_{n}$ divides its successor, ${\uparrow}^{C_{n}}_{B{\upharpoonright}_{n}}$ is total and bases are mapped to the next available multiple of the previous base, i.e., that $C_{n}$ is a good successor of $B{\upharpoonright}_{n}$, thus establishing 4.
  2. Case 2
    ($n\notin B$). In this case, we can only have $C_{n}\neq C_{n-1}$ if $n\in\partial B$ and $i>0$. Then, for $b={\rm Base}_{B}(n)$, we have that $1<b\mid n\notin B$ and $C_{n}=C_{n-1}\cup\{d_{j}\}_{j=1}^{i}$, where $d_{j}=d_{j}(n)$ are as given by Definition 5.1; namely, $d_{0}=c_{n-1}$ and $d_{j+1}=\langle\begin{smallmatrix}d_{j}\\ b\end{smallmatrix}\rangle n$. As observed previously, these values are finite due to the induction hypothesis applied to 4.
    Since $b\mid n$, $d_{j}\mid d_{j+1}$ by Lemma 3.81. Moreover, we cannot have $b=n$ since $n\notin B$, so $b<n$. Thus by Lemma 3.12, $d_{j}=\langle\begin{smallmatrix}d_{j}\\ b\end{smallmatrix}\rangle b<\langle\begin{smallmatrix}d_{j}\\ b\end{smallmatrix}\rangle n=d_{j+1}$. Thus, $C_{n}\setminus C_{n-1}=\{d_{j}\}_{j=1}^{n}$, with $c_{n-1}\mid d_{1}=S_{C_{n}}(c_{n-1})$ and $d_{j}\mid d_{j+1}=S_{C_{n}}(d_{j})$ for $1\leq j<i$, from which 1 follows.
@@ 2. Case 2
    If $x<n$, then ${\uparrow}x\leq{\uparrow}(n-1)$ by Lemma 3.11. Note that ${\rm Base}_{B}(n-1)=b$ since $n$ is not a base, so by the induction hypothesis applied to 3 and Lemma 3.12, ${\uparrow}(n-1)=\langle\begin{smallmatrix}c_{n-1}\\ b\end{smallmatrix}\rangle(n-1)<\langle\begin{smallmatrix}c_{n-1}\\ b\end{smallmatrix}\rangle n=d_{1}=S_{C_{n}}(c_{n-1})$. Item 2 then follows from Lemma 3.141.
@@
    According to Definition 2.5, ${\uparrow}_{n}n=\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle{\uparrow}_{n}n$, where $d\in C$ is least such that ${\uparrow}(n-1)<\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n<S_{C_{n}}(d)$, if such a $d$ exists. We claim that $d=d_{i}=c_{n}$. Since ${\uparrow}(n-1)<d_{1}<\langle\begin{smallmatrix}d_{i}\\ b\end{smallmatrix}\rangle n$ and $S_{C_{n}}(c_{n})=\infty$, it suffices to check that no $d<{c_{n}}$ is such that ${\uparrow}(n-1)<\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n<S_{C}(d)$. The induction hypothesis applied to 4 yields ${\uparrow}(n-1)=\langle\begin{smallmatrix}c_{n-1}\\ b\end{smallmatrix}\rangle(n-1)<\infty$. By Lemma 3.43, $d\geq c_{n-1}$. Thus $d\in C$ is such that $c_{n-1}\leq d<d_{i}$, so we must have that $d=d_{j}$ for some $j\in[0,i)$. However, $\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n=\langle\begin{smallmatrix}d_{j}\\ b\end{smallmatrix}\rangle n=d_{j+1}=S_{C}(d_{j})$, which rules out $d=d_{j}$, as this requires $\langle\begin{smallmatrix}d_{j}\\ b\end{smallmatrix}\rangle n<S_{C}(d_{j})$. We conclude that ${\uparrow}_{n}n=\langle\begin{smallmatrix}c_{n}\\ b\end{smallmatrix}\rangle n$, thus establishing 3. Item 4 then follows from the fact that $C_{n-1}$ is a good successor for $B{\upharpoonright}_{n}$, that $d_{j}\mid d_{j+1}=S_{C_{n}}(d_{j})$, and that ${\uparrow}_{n}n$ is well defined.
@@
With this, we are ready to prove that indeed $B\mapsto B_{+i}$ produces good successors.
@@
Let $B$ be a base hierarchy, $i\in\mathbb{N}$. Then, $B_{+i}$ is a good successor for $B$. In particular, $\mathcal{O}$ is a dynamical hierarchy.
Proof 5.8.
Let $_{+i}, =BC{\uparrow}={\uparrow}^{C}_{B}$, and $\langle\begin{smallmatrix}\cdot\\ \cdot\end{smallmatrix}\rangle=\langle\begin{smallmatrix}\cdot\\ \cdot\end{smallmatrix}\rangle^{C}_{B}$. For $n\in\mathbb{N}$, let $C_{n}=B_{+i}^{n}$, ${\uparrow}_{n}={\uparrow}^{C_{n}}_{B}$, and $\langle\begin{smallmatrix}\cdot\\ \cdot\end{smallmatrix}\rangle_{n}=\langle\begin{smallmatrix}\cdot\\ \cdot\end{smallmatrix}\rangle^{C_{n}}_{B}$. Then, $C$ is a base hierarchy because $\min C=\min B+1>2$, and if $c\in C$ and $S_{C}(c)<\infty$, then $c,S_{C}(c)\in C_{n}$ for large enough $n$, so $c\mid S_{C}(c)$ by Lemma 5.54.
To see that $C$ is a good successor for $B$, let $n\in\mathbb{N}$. If $S_{C}(n)<\infty$, let $m$ be large enough so that $S_{C}(n)\in C_{m}$, otherwise let $m=n$. We have that $C{\upharpoonright}_{S_{C}(n)}=C_{m}{\upharpoonright}_{S_{C}(n)}$ in view of Lemma 5.51 (as no bases smaller than $S_{C}(n)$ are added later), so Lemma 3.141 yields ${\uparrow}n={\uparrow}_{m}n<\infty$, so ${\uparrow}n$ is always finite. If moreover $n$ is a base, then ${\uparrow}_{m}n$ is the least multiple of ${\rm Base}_{B}(n-1)$ above ${\uparrow}_{m}(n-1)={\uparrow}(n-1)$, so $C$ satisfies the required condition on bases, and thus $C$ is a good successor for $B$.
It then follows from Theorem 2.16 that the fractal Goodstein process for $\mathcal{O}$ is always terminating; however, this fact is independent of $\sf KP$. In order to prove this, we need to review fundamental sequences for the Bachmann-Howard ordinal.
6 Fundamental Sequences

### v4

@@ 5. The Greedy Dynamical Hierarchy
We have shown that fractal Goodstein processes terminate, regardless of $\mathcal{B}$; in view of Example 2.15, this can be seen as an alternative proof of Goodstein’s original theorem if we set $\mathcal{B}_{i}=\{i+2\}$. However, for other choices of $\mathcal{B}$, the proof-theoretic strength of termination is much higher than that of the original theorem. In order to prove this, we will work with the greedy dynamical hierarchy of Definition 2.17. For the sake of generality, we also consider good successors of the form $B^{X}$ as given by Definition 2.10, where we defined $\partial B$ to be the set of $B$-critical numbers, i.e. the set of nonzero numbers that are multiples of their own base.
@@
Let $B$ be a base hierarchy and $i\in\mathbb{N}$. Let $X$ be such that $B\subseteq X\subseteq\partial B$, $C=B^{X}$, and for each $n\in\mathbb{N}$, $C_{n}=B^{X}_{n}$, $c_{n}=\max C_{n}$, ${\uparrow}_{n}={\uparrow}_{B}^{C_{n}}$, and $\langle\begin{smallmatrix}\cdot\\ \cdot\end{smallmatrix}\rangle_{n}=\langle\begin{smallmatrix}\cdot\\ \cdot\end{smallmatrix}\rangle_{B}^{C_{n}}$. Then, for every $n\in\mathbb{N}$ with $b={\rm Base}_{B}(n)$:
  1. (1)
@@
  4. (4)
    If $n\geq\min B$, then $C_{n}$ is a good successor for $B{\upharpoonright}_{n}$.
Proof.
@@
Proceed by induction on $n$. In this proof, ${\uparrow}={\uparrow}_{n-1}$ and $\langle\begin{smallmatrix}\cdot\\ \cdot\end{smallmatrix}\rangle=\langle\begin{smallmatrix}\cdot\\ \cdot\end{smallmatrix}\rangle_{n-1}$; note that ${\uparrow}x$ for $x<n$ and $\langle\begin{smallmatrix}y\\ x\end{smallmatrix}\rangle z$ when $2\leq x\leq n$ are finite by the induction hypothesis applied to (4) and Lemma 3.8.
@@
If $n=\min B$, then (4) is easy to check, since $C_{n}=\{\min B+1\}$ and $\min B+1>\min B\geq 2$, so $C_{n}$ is a good successor for $B{\upharpoonright}_{n}$. The other items are either also easy to check or hold vacuously. Similarly, if $n\notin X$, then $C_{n}=C_{n-1}$ and all claims follow directly from the induction hypothesis, so we may assume $n>\min B$ and $n\in X$, so that $C_{n}=C_{n-1}\cup\{c_{n}\}$, where $c_{n}$ is the least multiple of $c_{n-1}$ above ${\uparrow}(n-1)$. Let $b^{\prime}={\rm Base}_{B}(n-1)$.
@@
(1) By the induction hypothesis applied to (3) and the choice of $c_{n}$, $\langle\begin{smallmatrix}c_{n-1}\\ b^{\prime}\end{smallmatrix}\rangle(n-1)={\uparrow}_{n-1}(n-1)<c_{n}$, and since $\langle\begin{smallmatrix}c_{n-1}\\ b^{\prime}\end{smallmatrix}\rangle(n-1)\geq c_{n-1}$ by Lemma 3.1(4), we have that $c_{n}>c_{n-1}$.
(2) If $x<n$, then by Lemma 3.1(1), ${\uparrow}x\leq{\uparrow}(n-1)<c_{n}$, and since $C_{n-1}=C_{n}\upharpoonright_{c_{n}-1}$, (2) follows from Lemma 3.8(1).
@@
(3) According to Definition 2.5, ${\uparrow}_{n}n=\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n$, where $d\in C$ is least such that ${\uparrow}(n-1)<\langle\begin{smallmatrix}d\\ b\end{smallmatrix}\rangle n<S_{C_{n}}(d)$, if such a d{d} exists. We claim that $d=c_{n}$. If $n\in B$, this follows from Lemma 3.4(1), given that $c_{n}=S_{C_{n}}({\uparrow}(n-1))$. If $n\notin B$, since ${\uparrow}(n-1)<{c_{n}}<\langle\begin{smallmatrix}c_{n}\\ b\end{smallmatrix}\rangle n$ and $S_{C_{n}}(c_{n})=\infty$, it suffices to check that no ${d}<{c_{n}}$ with $d\in C_{n}$ is such that ${\uparrow}(n-1)<\langle\begin{smallmatrix}{d}\\ b\end{smallmatrix}\rangle n<S_{C}({d})$. The induction hypothesis applied to (4) yields ${\uparrow}(n-1)=\langle\begin{smallmatrix}c_{n-1}\\ b\end{smallmatrix}\rangle(n-1)<\infty$. By Lemma 3.3(3), ${d}\geq c_{n-1}$. Thus, ${d}\in\{c_{n-1},c_{n}\}$. So it remains to rule out $d=c_{n-1}$. However, since $n\in\partial B$, Lemma 3.5(1) yields $c_{n-1}\mid\langle\begin{smallmatrix}c_{n-1}\\ b\end{smallmatrix}\rangle n$. By the way we chose $c_{n}$, we conclude that $c_{n}\leq\langle\begin{smallmatrix}c_{n-1}\\ b\end{smallmatrix}\rangle n$, which rules out $d=c_{n-1}$, as this requires $\langle\begin{smallmatrix}c_{n-1}\\ b\end{smallmatrix}\rangle n<S_{C_{n}}(c_{n-1})$. We conclude that ${\uparrow}_{n}n=\langle\begin{smallmatrix}c_{n}\\ b\end{smallmatrix}\rangle n$.
(4) Since we chose $c_{n}$ to be the least multiple of $c_{n-1}$ above ${\uparrow}(n-1)$ and $C_{n-1}$ is a good successor for $B$ (and hence a base hierarchy) by the induction hypothesis, one readily checks using the previous items that every element of $C_{n}$ divides its successor, ${\uparrow}^{C_{n}}_{B{\upharpoonright}_{n}}$ is total, and bases are mapped to the next available multiple of the previous base; in other words, that $C_{n}$ is a good successor of $B{\upharpoonright}_{n}$. ∎
@@
With this, we are ready to prove that indeed $B\mapsto B^{X}$ produces good successors.
@@
Let $B$ be a base hierarchy, $B\subseteq X\subseteq\partial B$, $C=B^{X}$, and transformations be from $B$ to $C$.
  1. (1)
    $C$ is a good successor for $B$.
  2. (2)
    If $n\in X\setminus\{\min B\}$, then ${\rm Base}_{C}({\uparrow}(n-1))<{\rm Base}_{C}({\uparrow}n)$.
In particular, the above items hold for $\partial B$ in place of $X$ and $B^{+}$ in place of $C$, and the greedy dynamical hierarchy $\mathcal{D}$ of Definition 2.17 is a dynamical hierarchy.
Proof.
For $n\in\mathbb{N}$, let $C_{n}=B^{X}_{n}$, ${\uparrow}_{n}={\uparrow}^{C_{n}}_{B}$, and $\langle\begin{smallmatrix}\cdot\\ \cdot\end{smallmatrix}\rangle_{n}=\langle\begin{smallmatrix}\cdot\\ \cdot\end{smallmatrix}\rangle^{C_{n}}_{B}$.
(1) $C$ is a base hierarchy, because $\min C=\min B+1>2$, and if $c\in C$ and $S_{C}(c)<\infty$, then $c,S_{C}(c)\in C_{n}$ for large enough $n$, so $c\mid S_{C}(c)$ by Lemma 5.1(4).
To see that $C$ is a good successor for $B$, let $n\in\mathbb{N}$ and let $m$ be large enough so that $S_{C}(n)\in C_{m}$. We have that $C{\upharpoonright}_{S_{C}(n)}=C_{m}{\upharpoonright}_{S_{C}(n)}$ in view of Lemma 5.1(1) (as no bases smaller than $S_{C}(n)$ are added later), so Lemma 3.8(1) yields ${\uparrow}n={\uparrow}_{m}n<\infty$ and ${\uparrow}n$ is always finite. If moreover $n$ is a base, then ${\uparrow}_{m}n$ is the least multiple of ${\rm Base}_{B}(n-1)$ above ${\uparrow}_{m}(n-1)={\uparrow}(n-1)$, so $C$ satisfies the required condition on bases. We conclude that $C$ is a good successor for $B$.
(2) If $n\in X$, then by Lemma 5.1(3),
${\rm Base}_{C}({\uparrow}(n-1))=\max C_{n-1}<\max C_{n}={\rm Base}_{C}({\uparrow}n).\qed$
It then follows from Theorem 1 that the fractal Goodstein process for $\mathcal{D}$ is always terminating; however, this fact is independent of $\sf KP$. In order to prove this, we need to review fundamental sequences for the Bachmann-Howard ordinal.
6. Fundamental Sequences